OpenStudy (anonymous):

Find the smallest positive number x such that cosine squared ⁡ x minus 1.3cosine ⁡ x plus 0.42equals 0

OpenStudy (anonymous):

\[\cos ^2x-1.3\cos x+0.42\]

OpenStudy (anonymous):

i really dont know what to do

OpenStudy (anonymous):

@jim_thompson5910 @campbell_st Hello i was wondering if you can help me with this concept

jimthompson5910 (jim_thompson5910):

let z = cos(x) square both sides to get z^2 = cos^2(x)

jimthompson5910 (jim_thompson5910):

solve \[\Large z^2 - 1.3z + 0.42 = 0\] for z. Then use the solutions in terms of z to get the solutions in terms of x.